OpenStudy (anonymous):

If cos= 2 to the square root of 10 over 7 calculate the value of sin a. 3/7 b. square root of 29/7 c. square root of 69/7 d. square root of 89/7

OpenStudy (anonymous):

can you rewrite the question? as of now, it don't make no sense....

OpenStudy (anonymous):

\[\cos= 2\sqrt{10}/7\]

OpenStudy (campbell_st):

draw the triangle  find the missing side using pythagoras then sin = 3/7